OpenStudy (anonymous):

Pre-Calculus/Trigonometry Problem:

OpenStudy (anonymous):

Does \[\tan^{-1} (\tan \frac{ 2\pi }{ 3 })=\frac{ 2\pi }{ 3 }\] Why or why not?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

it's like \[f ^{-1} (f (x)) = x\]

OpenStudy (anonymous):

how would you explain that using 2pi/3

OpenStudy (anonymous):

my teacher wants us to explain it using that number

OpenStudy (anonymous):

it's the same idea. you just replace x with 2pi/3

OpenStudy (anonymous):

there must be some reasoning behind that

OpenStudy (campbell_st):

well the solution is partly correct... and tan is positive in 2 quadrants... so you need to look at the domain to see if there are any restictions before answering

OpenStudy (campbell_st):

oops 2pi/3 is negative.... but still 2 possible solutions withing the domain [0, 2pi]

OpenStudy (anonymous):

i don't know what to write as my answer... I'm kind of confused

OpenStudy (anonymous):

@Zale101 any ideas?

OpenStudy (zale101):

Let's say we have, \(sin(\theta)=0\) This equation interprets: The sine of what angle \(\theta\) is equal to ½ A good answer for this would be: "The angle whose sine is 0 is \(\pi\)" Which in an equation form, it would be written as this for the answer: \(sin^{-1}(\pi)=0\) \(tan^{-1}\) stands for tangent whose angle is \(tan(\frac{2\pi}{3})\) and the angle of tangent is in radian which is 2pi/3